# Non-Hermitian entanglement dip from scaling-induced exceptional criticality

### v1

@@ II.2 Effect of entanglement truncation interval in the effective 2-component model
The entanglement dip at $L\approx L_{c}$ occurs universally, regardless of the actual entanglement cut region $[x_{L},x_{R}]$, since it is not specifically dependent on the real-space profile of the matrix elements of $\bar{P}$. In Fig. S4(a) for our 2-component model Eq.S46, the odd-sized EE is shown to consistently exhibit $(L-L_{c})^{-0.5}$ divergence regardless of the truncation region.
This entanglement dip remains qualitatively similar (as it should be) in the EE of its parent 4-component coupled system [Eq. S1], even though the GBZ takes on a more intricate structure and the real part of the EP momentum is not rigorously fixed at $\pi$. Indeed, for this 4-component coupled system, the entanglement dip is observed in both odd and even-sized systems [See Fig. S4(b)], implying that it is not fundamentally based on the positions of the $\text{Re}K$ momentum points. What is fundamental is the fact that the EP lies on a GBZ whose imaginary part of momentum deviation $\Delta_{K}$ gives rise to a $(L-L_{c})^{-0.5}$ divergence in $P$.
@@ II.2.1 Single unit-cell truncation
Since the entanglement dip is largely independent of the truncation region, we can make further analytic progress by focusing on single-site entanglement where only one unit-cell $x_{L}=x_{R}$ exists in the untruncated region $[x_{L},x_{R}]=[x_{L},x_{L}]$. For a single unit-cell, the Fourier phase factor $\text{e}^{ik(x_{1}-x_{2})}$ in $\langle{x,\alpha}|P|x,\beta\rangle$ [Eq. S59] disappears, and the single unit-cell truncated projector (we call it $\tilde{P}$ here to emphasize the single unit cell special case) is
@@
$\begin{split}&\tilde{p}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{Approx.}}}=\frac{1}{2}\pm\sqrt{\tilde{P}^{+-}\times\tilde{P}^{-+}}\\ &=\frac{1}{2}\pm\left(\displaystyle\frac{1}{4}{\displaystyle\frac{t_{R}}{t_{L}}}\sqrt{\displaystyle\frac{t_{L}/t_{R}-1/t^{2}_{R}}{\alpha}}\frac{1}{\sqrt{L-L_{c}}}+\frac{1}{2\pi\sqrt{t_{R}t_{L}}}\sqrt{(t_{L}+1)(t_{R}+1)}\right)+\mathcal{O}\ ,\end{split}$ (S65)
@@
where $a=\frac{1}{2\pi\sqrt{t_{R}t_{L}}}\sqrt{(t_{L}+1)(t_{R}+1)}+\mathcal{O}$, $b=\frac{1}{4}{\frac{t_{R}}{t_{L}}}\sqrt{\frac{t_{L}/t_{R}-1/t^{2}_{R}}{\alpha}}$, $\mathcal{O}$ and $\mathcal{O}^{\prime}$ representing small offsets that can be adjusted to mitigate the imperfect approximation [See Fig. S6(b)].
Figure S6: Scaling behavior of the $\bar{P}$ spectrum and the EE of the 2-component model [Eq. S46] with single unit-cell truncation. (a) The excellent agreement between the numerical spectra of $\bar{P}$ [Eq. LABEL:2bandP3] and its single unit-cell approximation $\tilde{P}$ [Eq. S63], as well as its analytically approximated spectrum $\tilde{p}_{\text{Approx.}}$ [Eq. S65]. Note that $\tilde{P}$ only yields the nontrivial pair of eigenvalues away from 0 and 1, whose kink near $L\approx L_{c}$ is well-approximated by $\tilde{p}_{\text{Approx.}}$. (b) The corresponding scaling of the entanglement entropy (EE) $S(\bar{P})$ and $S(\tilde{P})$ from the eigenvalues of $\bar{P}$ and $\tilde{P}$ shown in panel (a), along with the analytical result $S_{\text{Approx.}}$ given by Eq. S66. The results exhibit good consistency, and in particular all predict the same entanglement dip. Parameters are $t_{L}=1.2\text{e}^{0.3},t_{R}=1.2\text{e}^{-0.3},\delta=1.6\times 10^{-3}$. The coefficients $a$ and $b$ in Eq. S66 are calculated as $a=0.3905$ and $b=0.0456$ with small constants $\mathcal{O}=0.14$, $\mathcal{O}^{\prime}=-0.035$ [Eq.S65,S66], corresponding to the critical scale being $L_{c}=42.79$.
@@ II.2.2 Effect of the length of the entanglement truncation interval
In Hermitian systems, it is well-known that the EE of a cut region $[1,n_{cut}]$ varies strongly like $S\sim\log\left(L\sin\frac{\pi n_{cut}}{L}\right)$, as can be proven with boundary CFTcite. Hence it is prudent to check whether SIEC behavior is nontrivially influenced by $n_{cut}$.
@@ II.3 Controlling the depth of entanglement dips
Below, we explore the relationship between the depth of the entanglement dip (i.e. minumum EE $S_{\text{min}}$) varying the coupling parameter $\delta$ while keeping other parameters constant. This approach is based on our analytical understanding of how $L_{c}$ varies with $\delta$ [See Fig. S2]. Here, we numerically identify parameters that lead to exceptionally negative $S_{\text{min}}$ values, where the entanglement dip represents drastic departures from usual $\log L$ entanglement scaling.
Although the theoretically predicted $L_{c}$ varies continuously with $\delta$ [Eq. LABEL:Eqcritical1], an actual lattice contains only an integer number of unit cells $L$. This prevents us from getting infinitesimally close to $L_{c}$, where $S$ truly diverges. Away from that, how negative a dip $S$ can reach depends on commensurability considerations, as shown in Fig. S8. For certain fine-tuned values of $\delta$, the EE through $S_{\text{min}}$ can dip below $-4$ (such as at $\delta$=1.68441635 ×$\times 10^{-3}$ or $2.131947\times 10^{-3}$), corresponding to extremely large $p$ eigenvalues. Such dips occur both for our effective two-component model (a) as well as its parent four-component model (b), even though the dip positions are different. But what remains consistent are the shapes and qualitative order of magnitude of the entanglement dips.
Figure S8: Sensitivity of the entanglement dip minimum $S_{\text{min}}$ with coupling $\delta$. The entanglement dip $S_{\text{min}}$ can yield negative EE for a wide range of parameters, and can even reach very negative values of $S_{\text{min}}<-4$ across small parameter windows. Shown are the $S_{\text{min}}$ and the corresponding system chain length $L$ where it is attained, both for our (a) 2-component effective model Eq. S46 with $\delta$ built into its scale-dependent GBZ and (b) its parent 4-component model Eq. S1 with $\delta$ being the physical inter-chain coupling. A more complete parameter space plot that shows the dependence of $S_{\text{min}}$ with $t_{L}/t_{R}$ is given in Fig. 3(c) of the main text. Parameters are $t_{L}=1.2\text{e}^{0.3},t_{R}=1.2\text{e}^{-0.3}$.
@@ II.4 Suppressed NHSE in the entanglement eigenstates
But, as evident from Figs. S9(c) and S10(c), the entanglement eigenstate $\Psi_{p_{m}}$ shows a weaker skin localization than $\Psi_{\text{EP}}$, despite being approximately made up of skin modes. This can be interpreted as a signature of NHSE suppression in the entanglement eigenstate, and can be qualitatively understood as a consequence of the antagonistic competition between the oppositely directed NHSE chains.
Figure S9: Entanglement scaling and $\bar{P}$ eigenstate profile for our physical 4-component model Eq. S1. (a) $\bar{P}$ spectrum as a function of system size $L$ at half entanglement truncation. A conventional in-gap spectral branch for critical systems appears for $L>L_{c}$. (b) The resultant entanglement entropy scaling, with a pronounced entanglement dip before $L\approx L_{c}$. (c1) The spatial profiles of $\Psi_{p_{m}}$(the $S=S_{\text{min}}$ eigenstate of $\bar{P}$, colored red above) and $\Psi_{\text{EP}}$ (eigenstate of Hamiltonian $\mathcal{H}$) are shown. Since $\bar{P}$ is constructed using the eigenstates of the Hamiltonian, its eigenstate $\Psi_{p_{m}}$ inherits the skin effect from the Hamiltonian. (c2) To eliminate the skin effect in $\Psi_{p_{m}}$ and $\Psi_{\text{EP}}$, the profile of $|\Psi_{p_{m}}/\Psi_{\text{EP}}|$ represents the cumulative wavefunction distribution near the truncation point after excluding the EP eigenstate. Parameters used are $t_{L}=1.2\text{e}^{0.3},t_{R}=1.2\text{e}^{-0.3}$, ${\delta=1.6\times 10^{-3}}$.
Figure S10: Entanglement scaling and $\bar{P}$ eigenstate profile for our physical 4-component model Eq. S1, for an extremely pronounced entanglement dip. This figure is similar to Fig. S9, but with inter-chain coupling tuned to $\delta=1.6844163\times 10^{-3}$ such that $L_{c}=40.038$ becomes almost an integer, giving rise to very negative $S_{\text{min}}$. (a) and (b) shows how the $\bar{P}$ spectrum and corresponding entanglement entropy jump dramatically at $L\approx L_{c}$. However, as compared to Fig. S9(c1,c2), the spatial distribution of $\Psi_{p_{m}}$ (depicted in (c1) and colored red in (a,b)) and $|\Psi_{p_{m}}/\Psi_{\text{EP}}|$ in (c2) do not exhibit obvious changes. Parameters are $t_{L}=1.2\text{e}^{0.3},t_{R}=1.2\text{e}^{-0.3}$.
@@ III Comparison with other types of critical scenarios
IV Measuring Negative biorthogonal entanglement
@@
where $K$ is complex conjugation operation and $|\Psi^{R}_{E}\rangle$ and $|\Psi^{L}_{E}\rangle$ represent the eigenstates of the Hamiltonian $\mathcal{H}$ and its hermitian conjugate $\mathcal{H}^{\dagger}$ also with eigenvalue $E$.

### v2

@@ II.2 Effect of entanglement truncation interval in the effective 2-component model
The entanglement dip at $L\approx L_{c}$ occurs universally, regardless of the actual entanglement cut region $[x_{L},x_{R}]$, since it is not specifically dependent on the real-space profile of the matrix elements of $\bar{P}$. In Fig. S4a for our 2-component model Eq. (S46), the odd-sized EE is shown to consistently exhibit $(L-L_{c})^{-0.5}$ divergence regardless of the truncation region.
This entanglement dip remains qualitatively similar (as it should be) in the EE of its parent 4-component coupled system (Eq. (S1)), even though the GBZ takes on a more intricate structure and the real part of the EP momentum is not rigorously fixed at $\pi$. Indeed, for this 4-component coupled system, the entanglement dip is observed in both odd and even-sized systems (See Fig. S4b), implying that it is not fundamentally based on the positions of the $\text{Re}K$ momentum points. What is fundamental is the fact that the EP lies on a GBZ whose imaginary part of momentum deviation $\Delta_{K}$ gives rise to a $(L-L_{c})^{-0.5}$ divergence in $P$.
@@ II.2.1 Single unit-cell truncation
Since the entanglement dip is largely independent of the truncation region, we can make further analytic progress by focusing on single-site entanglement where only one unit-cell $x_{L}=x_{R}$ exists in the untruncated region $[x_{L},x_{R}]=[x_{L},x_{L}]$. For a single unit-cell, the Fourier phase factor $\text{e}^{ik(x_{1}-x_{2})}$ in $\langle{x,\alpha}|P|x,\beta\rangle$ (Eq. (S59)) disappears, and the single unit-cell truncated projector (we call it $\tilde{P}$ here to emphasize the single unit cell special case) is
@@
$\begin{split}&\tilde{p}_{{\text{Approx.}}}=\frac{1}{2}\pm\sqrt{\tilde{P}^{+-}\times\tilde{P}^{-+}}\\ &=\frac{1}{2}\pm\left(\displaystyle\frac{1}{4}{\displaystyle\frac{t_{R}}{t_{L}}}\sqrt{\displaystyle\frac{t_{L}/t_{R}-1/t^{2}_{R}}{\alpha}}\frac{1}{\sqrt{L-L_{c}}}+\frac{1}{2\pi\sqrt{t_{R}t_{L}}}\sqrt{(t_{L}+1)(t_{R}+1)}\right)+\mathcal{O}\ ,\end{split}$ (S65)
@@
where $a=\frac{1}{2\pi\sqrt{t_{R}t_{L}}}\sqrt{(t_{L}+1)(t_{R}+1)}+\mathcal{O}$, $b=\frac{1}{4}{\frac{t_{R}}{t_{L}}}\sqrt{\frac{t_{L}/t_{R}-1/t^{2}_{R}}{\alpha}}$, $\mathcal{O}$ and $\mathcal{O}^{\prime}$ representing small offsets that can be adjusted to mitigate the imperfect approximation (See Fig. S6b).
Figure S6: Scaling behavior of the $\bar{P}$ spectrum and the EE of the 2-component model (Eq. (S46)) with single unit-cell truncation. (a) The excellent agreement between the numerical spectra of $\bar{P}$ (Eq. (LABEL:2bandP3)) and its single unit-cell approximation $\tilde{P}$ (Eq. (S63)), as well as its analytically approximated spectrum $\tilde{p}_{\text{Approx.}}$ (Eq. (S65)). Note that $\tilde{P}$ only yields the nontrivial pair of eigenvalues away from 0 and 1, whose kink near $L\approx L_{c}$ is well-approximated by $\tilde{p}_{\text{Approx.}}$. (b) The corresponding scaling of the entanglement entropy (EE) $S(\bar{P})$ and $S(\tilde{P})$ from the eigenvalues of $\bar{P}$ and $\tilde{P}$ shown in panel (a), along with the analytical result $S_{\text{Approx.}}$ given by Eq. (S66). The results exhibit good consistency, and in particular all predict the same entanglement dip. Parameters are $t_{L}=1.2\text{e}^{0.3},t_{R}=1.2\text{e}^{-0.3},\delta=1.6\times 10^{-3}$. The coefficients $a$ and $b$ in Eq. (S66) are calculated as $a=0.3905$ and $b=0.0456$ with small constants $\mathcal{O}=0.14$, $\mathcal{O}^{\prime}=-0.035$ (Eq. (S65,S66)), corresponding to the critical scale being $L_{c}=42.79$.
@@ II.2.2 Effect of the length of the entanglement truncation interval
In Hermitian systems, it is well-known that the EE of a cut region $[1,n_{cut}]$ varies strongly like $S\sim\log\left(L\sin\frac{\pi n_{cut}}{L}\right)$, as can be proven with boundary CFT. Hence it is prudent to check whether SIEC behavior is nontrivially influenced by $n_{cut}$.
@@ II.3 Controlling the depth of entanglement dips
Below, we explore the relationship between the depth of the entanglement dip (i.e. minumum EE $S_{\text{min}}$) varying the coupling parameter $\delta$ while keeping other parameters constant. This approach is based on our analytical understanding of how $L_{c}$ varies with $\delta$ (See Fig. S2). Here, we numerically identify parameters that lead to exceptionally negative $S_{\text{min}}$ values, where the entanglement dip represents drastic departures from usual $\log L$ entanglement scaling.
Although the theoretically predicted $L_{c}$ varies continuously with $\delta$ (Eq. (LABEL:Eqcritical1)), an actual lattice contains only an integer number of unit cells $L$. This prevents us from getting infinitesimally close to $L_{c}$, where $S$ truly diverges. Away from that, how negative a dip $S$ can reach depends on commensurability considerations, as shown in Fig. S8. For certain fine-tuned values of $\delta$, the EE through $S_{\text{min}}$ can dip below $-4$ (such as at $\delta$=1.68441635 ×$\times 10^{-3}$ or $2.131947\times 10^{-3}$), corresponding to extremely large $p$ eigenvalues. Such dips occur both for our effective two-component model (a) as well as its parent four-component model (b), even though the dip positions are different. But what remains consistent are the shapes and qualitative order of magnitude of the entanglement dips.
Figure S8: Sensitivity of the entanglement dip minimum $S_{\text{min}}$ with coupling $\delta$. The entanglement dip $S_{\text{min}}$ can yield negative EE for a wide range of parameters, and can even reach very negative values of $S_{\text{min}}<-4$ across small parameter windows. Shown are the $S_{\text{min}}$ and the corresponding system chain length $L$ where it is attained, both for our (a) 2-component effective model Eq. (S46) with $\delta$ built into its scale-dependent GBZ and (b) its parent 4-component model Eq. (S1) with $\delta$ being the physical inter-chain coupling. A more complete parameter space plot that shows the dependence of $S_{\text{min}}$ with $t_{L}/t_{R}$ is given in Fig. 3(c) of the main text. Parameters are $t_{L}=1.2\text{e}^{0.3},t_{R}=1.2\text{e}^{-0.3}$.
@@ II.4 Suppressed NHSE in the entanglement eigenstates
But, as evident from Figs. S9c and S10c, the entanglement eigenstate $\Psi_{p_{m}}$ shows a weaker skin localization than $\Psi_{\text{EP}}$, despite being approximately made up of skin modes. This can be interpreted as a signature of NHSE suppression in the entanglement eigenstate, and can be qualitatively understood as a consequence of the antagonistic competition between the oppositely directed NHSE chains.
Figure S9: Entanglement scaling and $\bar{P}$ eigenstate profile for our physical 4-component model Eq. (S1). (a) $\bar{P}$ spectrum as a function of system size $L$ at half entanglement truncation. A conventional in-gap spectral branch for critical systems appears for $L>L_{c}$. (b) The resultant entanglement entropy scaling, with a pronounced entanglement dip before $L\approx L_{c}$. (c1) The spatial profiles of $\Psi_{p_{m}}$(the $S=S_{\text{min}}$ eigenstate of $\bar{P}$, colored red above) and $\Psi_{\text{EP}}$ (eigenstate of Hamiltonian $\mathcal{H}$) are shown. Since $\bar{P}$ is constructed using the eigenstates of the Hamiltonian, its eigenstate $\Psi_{p_{m}}$ inherits the skin effect from the Hamiltonian. (c2) To eliminate the skin effect in $\Psi_{p_{m}}$ and $\Psi_{\text{EP}}$, the profile of $|\Psi_{p_{m}}/\Psi_{\text{EP}}|$ represents the cumulative wavefunction distribution near the truncation point after excluding the EP eigenstate. Parameters used are $t_{L}=1.2\text{e}^{0.3},t_{R}=1.2\text{e}^{-0.3}$, ${\delta=1.6\times 10^{-3}}$.
Figure S10: Entanglement scaling and $\bar{P}$ eigenstate profile for our physical 4-component model Eq. (S1), for an extremely pronounced entanglement dip. This figure is similar to Fig. S9, but with inter-chain coupling tuned to $\delta=1.6844163\times 10^{-3}$ such that $L_{c}=40.038$ becomes almost an integer, giving rise to very negative $S_{\text{min}}$. (a) and (b) shows how the $\bar{P}$ spectrum and corresponding entanglement entropy jump dramatically at $L\approx L_{c}$. However, as compared to Fig. S9c1 and c2, the spatial distribution of $\Psi_{p_{m}}$ (depicted in (c1) and colored red in (a,b)) and $|\Psi_{p_{m}}/\Psi_{\text{EP}}|$ in (c2) do not exhibit obvious changes. Parameters are $t_{L}=1.2\text{e}^{0.3},t_{R}=1.2\text{e}^{-0.3}$.
@@ III Comparison with other types of critical scenarios
IV Generality of entanglement dips
SIEC and entanglement dips are not limited to the coupled $H_{\text{SSH}}$ chains explicitly computed so far. They are expected to show up whenever the spectrum depends on $L$, and becomes gapless at an EP at a special $L\approx L_{c}$. This requires a scale-dependent GBZ that can be generically designed by coupling subsystems with competing NHSE pumping. Presented in Fig. S13 are other models containing oppositely-directed NHSE channels, albeit with lower symmetry. Evidently, they all exhibit scale-dependent GBZs, and importantly exhibit entanglement dips at $L\approx L_{c}$ whenever the GBZ encounters a critical point $K_{0}$ (red).
Figure S13: Entanglement dips and scale-dependent GBZs in generalized models satisfying Eq. (S1). (a–c) $L$-dependent generalized Brillouin zones (GBZs) for 3 illustrative models: (a) $H=(t_{L}+1/z)\sigma_{+}+(t_{R}+z)\sigma_{-}$, (b) $H=(t_{L}+\frac{1}{3}z^{-2})\sigma_{+}+(t_{R}+z)\sigma_{-}$, and (c) $H=t_{L}\sigma_{+}+(0.6+0.12i+z+0.1z^{-2})\sigma_{-}$. All models use weak inter-chain coupling $\delta=1.6\times 10^{-3}$ and asymmetric hopping $t_{L},t_{R}=1.2e^{\pm 0.3}$. For each case, GBZs are shown as a function of system size $L$, increasing from top to bottom. (d) Entanglement entropy $S$ exhibits clear dips at critical system sizes where the GBZ loops intersect the special momentum point $K_{0}$ (highlighted in red), despite the GBZs having irregular shapes.
V Measuring Negative biorthogonal entanglement
@@
where $K$ is complex conjugation operation and $|\Psi^{R}_{E}\rangle$ and $|\Psi^{L}_{E}\rangle$ represent the eigenstates of the Hamiltonian $\mathcal{H}$ and its hermitian conjugate $\mathcal{H}^{\dagger}$ ( $\mathcal{H}^{\dagger}=\mathcal{H}^{\text{T}}$) also with eigenvalue $E$.
@@
Additionally, from an application perspective, active photonic systems have emerged as promising platforms for realizing non-Hermitian effects, including the skin effect studied here. In particular, hybrid photonic setups such as exciton-polariton condensates have demonstrated controllable gain and loss mechanisms that naturally implement non-Hermitian Hamiltonians [37, 38]. These systems allow for precise engineering of boundary conditions and coupling strengths, enabling the observation of non-Hermitian skin effects and related topological phenomena in a highly tunable environment. Furthermore, the inherent nonlinearity and driven-dissipative nature of active photonic platforms offer unique opportunities for dynamic control and switching applications in photonic devices. Therefore, the measurement schemes and theoretical frameworks discussed here could be adapted to explore and harness these effects in active photonic devices, enhancing their potential for technological applications.
Moreover, the simulation of non-Hermitian systems on quantum computing platforms has achieved remarkable progress [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. Specifically, the biorthogonal measurements required for a non-Hermitian system $\mathcal{H}$ can be reformulated as joint measurements of an auxiliary operator $\hat{\Lambda}$ and a physical observable $A$ on the right eigenstate of a larger system $\tilde{\mathcal{H}}$(Eq. (S67)), as shown in Eq. (S70). In quantum computing experiments, even the measurement of non-unitary operators can be efficiently implemented using advanced quantum algorithms and error mitigation techniques. This approach not only significantly broadens the experimental toolkit for validating non-Hermitian phenomena but also provides a solid technical foundation for exploring and harnessing unconventional quantum mechanical effects.